kim was charged 560.00 interest for 1 year on a 10,000.00 credit card balance what was her monthly interest rate please show work
Answer by mananth(12270)

interest = 560
credit card balance = 10000
...
I=p*r*t
560=10000*r*1
560/10000= r
r=0.56
0.56*100 = 5.6% for 1 year.
monthly rate was 5.6 /12 =0.46 % per month
...
m.ananth@hotmail.ca